Circular No. 8850

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V390 NORMAE = NOVA NORMAE 2007 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag ≈ 9.4) located at α = 16^{h}32^{m}.2, δ = −45^{o}08′ (equinox 2000.0) on two Technical Pan films taken around June 15.086 UT with an orange filter and an 85-mm camera lens. G. Bolt (Perth, Western Australia, 0.20-m reflector) reports that his unfiltered CCD image taken on June 16.7 yields the following precise position for the new object: α = 16^{h}32^{m}11^{s}.51, δ = −45^{o}09′13′′.4. Additional reported CCD position end figures: 11^{s}.46, 13′′.6 (R. Santallo, Tahiti, 0.20-m reflector); 11^{s}.48, 13′′.2 (L. A. G. Monard, Pretoria, S. Africa, 0.35-m reflector). Additional approximate magnitudes: 1997 Mar. 31, [20 (Digitized Sky Survey, via Bolt); 2007 May 17.615, [10.2 (K. Haseda, Toyohashi, Aichi, Japan, Canon EOS-5D camera + 120-mm f/3.5 telephoto lens; via H. Yamaoka, Kyushu University); 20.1, [11.0 (Liller); 27.701, 9.3 (Haseda; prediscovery); June 7.193, 9.1 (Liller); 16.7, 8.9-9.5 (Bolt); 16.946, 10.3 (J. G. de S. Aguiar, Campinas, Brazil; visual); 17.76, B = 11.0, V = 10.1, R = 9.3 (Monard); 17.950, 10.1 (Aguiar; visual); 18.042, 10.3 (A. Amorim, Florianopolis, Brazil; visual); 18.498, 9.3 (Santallo; CCD red); 18.520, 10.2 (A. Pearce, Nedlands, W. Australia; visual); 18.945, 10.0 (Aguiar; visual). Bolt notes that the nearest star to the new object appears to be a USNO-B1.0-catalogued star (red mag ∼ 15–16) located ∼ 3′′ to the northwest.

E. Kazarovets, Sternberg Astronomical Institute, reports that the designation V390 Nor has been assigned to this nova.

COMETS C/2004 A4, C/2004 B11, AND C/2004 J20 (SOHO) Three additional near-sun comets (cf. IAUC 8848) have been found

by B. Zhou on archival 2004 SOHO website images — all being Kreutz sungrazers except for C/2004 J20 (Kracht group). Each was stellar in appearance and faint, the two Kreutz sungrazings reaching mag ∼ 7–7.5 and C/2004 J20 peaking in brightness at mag ∼ 8.

Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 A4 Jan. 7.929 19^{h}25.^{m}6 −25^{o}43′ C3 BZ 2007-L16
C/2004 B11  20.864 20 22.6 −21 54 C3 BZ 2007-L16
C/2004 J20 May 13.618 3 27.2 +19 11 C2 BZ 2007-K65
2007 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green